# Detect duality obstruction of calibrations in smooth category

### v2

@@ 6.2. Generalization to higher codimension
It is clear that those in (6.2) are in the same homological class by our construction. Let us explain why they are homologically mass-minimizing. By Nash’s embedding theorem [Nash], we have isometric embedding $F_{i}:(X_{i},\tilde{h}_{i})\longrightarrow(\mathbb{R}^{\ell_{i}},g_{E})$ and then one can pull back $\tilde{\Phi}_{i}$ to a form $\bar{\Phi}_{i}=\pi_{i}^{*}(\tilde{\Phi}_{i})$ in a small neighborhood of $F_{i}(X_{i})$, where $\pi_{i}$ is the nearest projection to $F_{i}(X_{i})$. Now we can perform mollifications to get a series of smooth forms $\big(\bar{\Phi}_{i}\big)_{\epsilon}$ in a smaller neighborhood $U$ of $F_{i}(X_{i})$ and restrict them back to get smooth closed forms $\varphi_{i}^{\epsilon}$ in $F_{i}(X_{i})$.
It should be pointed out that the mollification through singular set can be done due to the following. As we consider $\varphi=d(h_{2}\psi)$ for Lawlor cone where $h_{2}$ is from §4.2 and our extension of $\varphi$ to $\Phi$ on neighborhood $\Xi$ of $S$ by (3.4) shows that $\Phi$ is exact (and simple, with a smooth bounded potential $\omega$) away from the singular points in $\Xi$. Hence, one can apply the mollification to $\bar{\Phi}_{i}$ to have smooth forms $\big(\bar{\Phi}_{i}\big)_{\epsilon}$ in the Euclidean space $(\mathbb{R}^{\ell_{i}},g_{E})$. Moreover, as $\tilde{\Phi}_{i}$ and $\Phi_{i}$ are the same around $S_{i}$, in particular around the set $\pi_{i}^{-1}S_{i}\cap U$ we have
(6.3) $d\big(\bar{\Phi}_{i}\big)_{\epsilon}=\big(d\bar{\Phi}_{i}\big)_{\epsilon}=\big(\pi_{i}^{*}(dd\omega_{i})\big)_{\epsilon}=0.$
Note that mollification can be done for forms whenever they are defined almost everywhere, locally bounded and measurable. In the above we actually apply mollification for the $L^{1}_{loc}$-form $\pi_{i}^{*}\omega_{i}$ to have $d\big(\pi_{i}^{*}\omega_{i}\big)_{\epsilon}=\big(\bar{\Phi}_{i}\big)_{\epsilon}$ which is stronger than (6.3) in the local.
By (6.3) we know that $d\varphi_{i}^{\epsilon}=0$. So we have the equal integrals for terms in (6.2)
@@
According to the comass control method in the proof of Lemma 6.3, we know that the comass $\|\varphi^{\epsilon}_{1}\wedge\varphi^{\epsilon}_{2}\|^{*}_{\tilde{h}_{1}\oplus\tilde{h}_{2}}<\frac{1}{(m_{1}+m_{2})!}$ in $U^{c}_{1}\times U^{c}_{2}$ and the sequence $\{\varphi^{\epsilon}_{1}\wedge\varphi^{\epsilon}_{2}\}_{\epsilon\downarrow 0}$ serves as a calibration in the sense that, for any compactly supported competitor $T$ of normal current in $M_{1}\times M_{2}]$,
$\mathrm{\mathbf{M}}(M_{1}\times M_{2})=\lim_{\epsilon\downarrow 0}\int_{M_{1}\times M_{2}}\varphi_{1}^{\epsilon}\wedge\varphi_{2}^{\epsilon}=\lim_{\epsilon\downarrow 0}\int_{T}\varphi_{1}^{\epsilon}\wedge\varphi_{2}^{\epsilon}\leq\mathrm{\mathbf{M}}(T)$
@@
$\lim_{\epsilon\downarrow 0}\max_{X_{1}\times X_{2}}\|\varphi^{\epsilon}_{1}\wedge\varphi^{\epsilon}_{2}\|^{*}_{\tilde{h}_{1}\oplus\tilde{h}_{2}}=1$
@@
We can repeat the argument before Lemma 6.3 for contradiction. Fix $q\in M_{1}$ and $p=p_{1,2}$ a singular point of $S_{2}$ and set $x=(q,p)$. Let $\xi$ be the oriented unit top vector of $T_{q}M_{1}$. Then $\xi\llcorner\varpi_{x}$ is a calibration form with constant coefficients in $\left(T_{x}(X_{1}\times X_{2}),(\tilde{h}_{1}\oplus\tilde{h}_{2})_{x}\right)$. Since $\varpi$ calibrates $M_{1}\times S_{2}$, it can be seen that the restriction of $\xi\llcorner\varpi_{x}$ to $\{0\}\times\mathbb{R}^{N_{2}+1}$ calibrates the tangent cone of $S_{2}$ at $p$. This contradicts with our construction which is based on local model from Theorem 1.4 - the tangent cone of $S_{2}$ at each singular point cannot support any calibration with constant coefficients.

### v3

@@ 6.2. Generalization to higher codimension
It is clear that those in (6.2) are in the same homological class by our construction. Let us explain why they are homologically mass-minimizing. By Nash’s embedding theorem [Nash], we have isometric embedding $F_{i}:(X_{i},\tilde{h}_{i})\longrightarrow(\mathbb{R}^{\ell_{i}},g_{E})$ and then one can pull back $\tilde{\Phi}_{i}$ to a form $\bar{\Phi}_{i}=\pi_{i}^{*}(\tilde{\Phi}_{i})$ in a small neighborhood of $F_{i}(X_{i})$, where $\pi_{i}$ is the nearest projection to $F_{i}(X_{i})$. Now we can perform mollifications to get a family of smooth forms $\big(\bar{\Phi}_{i}\big)_{\epsilon}$ in a small neighborhood of $F_{i}(X_{i})$ and restrict them back to get smooth closed forms $\varphi_{i}^{\epsilon}$ in $F_{i}(X_{i})$.
Mollification through singular set can be done due to the following. By §4.2 and §3.2 the resulting $\tilde{\Phi}_{i}=d\omega_{i}$ around $o=p_{i,1}$ (or $p_{i,2}$) for $\omega_{i}=h_{2}\psi_{i}$ (smooth and of bounded derivatives away from $o$ in local). Therefore,  [^3] Cf. the proof of Corollary 6.5 for more details.
(6.3) $\big(\bar{\Phi}_{i}\big)_{\epsilon}=\big(\pi_{i}^{*}d\omega_{i}\big)_{\epsilon}=d\big(\pi_{i}^{*}\omega_{i}\big)_{\epsilon}\text{\ \ \ \ \ \ and \ \ \ \ \ \ }d\big(\bar{\Phi}_{i}\big)_{\epsilon}=0,$
so not hard to see that the restriction $\varphi_{i}^{\epsilon}$ of smooth $\big(\bar{\Phi}_{i}\big)_{\epsilon}$ is also closed on the entire $F_{i}(X_{i})$ (identified as $X_{i}$ through the embedding).
By (6.3) we know that $d\varphi_{i}^{\epsilon}=0$ and thus $d\big(\varphi_{1}^{\epsilon}\wedge\varphi_{2}^{\epsilon}\big)=0$. So we have the equal integrals for terms in (6.2)
@@
According to the comass control method in the proof of Lemma 6.3, we know that the comass $\|\varphi^{\epsilon}_{1}\wedge\varphi^{\epsilon}_{2}\|^{*}_{\tilde{h}_{1}\oplus\tilde{h}_{2}}<\frac{1}{(m_{1}+m_{2})!}$ in $U^{c}_{1}\times U^{c}_{2}$ when $\epsilon$ is small and the sequence $\{\varphi^{\epsilon}_{1}\wedge\varphi^{\epsilon}_{2}\}_{\epsilon\downarrow 0}$ serves as a calibration in the sense that, for any compactly supported competitor $T$ of normal current in $M_{1}\times M_{2}]$,
(6.5) $\mathrm{\mathbf{M}}(M_{1}\times M_{2})=\lim_{\epsilon\downarrow 0}\int_{M_{1}\times M_{2}}\varphi_{1}^{\epsilon}\wedge\varphi_{2}^{\epsilon}=\lim_{\epsilon\downarrow 0}\int_{T}\varphi_{1}^{\epsilon}\wedge\varphi_{2}^{\epsilon}\leq\mathrm{\mathbf{M}}(T)$
@@
$\lim_{\epsilon\downarrow 0}\max_{X_{1}\times X_{2}}\|\varphi^{\epsilon}_{1}\wedge\varphi^{\epsilon}_{2}\|^{*}_{\tilde{h}_{1}\oplus\tilde{h}_{2}}=1.$
@@
A corollary of the above is the minimality of the Cartesian product of two Lawlor cones. Here we provide a direct proof in the Euclidean space.
Corollary 6.5.
Let $C_{1}\subset\mathbb{R}^{N_{1}+1}$ and $C_{2}\subset\mathbb{R}^{N_{2}+1}$ be two Lawlor cones. Then the cone $C_{1}\times C_{2}\subset\mathbb{R}^{N_{1}+1}\oplus\mathbb{R}^{N_{2}+1}$ (not a regular cone) is also mass minimizing.
Proof.
Let $d\omega_{i}$ be the coflat calibration constructed in §4.2 for $C_{i}$. Note that $\omega_{i}$ is singular (not defined) only at the origin, and smooth away from the origin. Take $\sigma$ to be a mollifier, which is a smooth function supported in the unit ball of $\mathbb{R}^{n}$ centered at the origin with total integral one. Set $\sigma_{\epsilon}(z)=\frac{1}{\epsilon^{n}}\sigma(\epsilon z)$. Then mollification
$f_{\epsilon}(x)=\int_{\mathbb{R}^{n}}f(y)\sigma_{\epsilon}(y-x)dy$
is well defined for almost everywhere defined real-valued, locally bounded, measurable function $f$. By mollifying $\omega_{i}$ we mean the mollification on its coefficient functions $f_{i,J}$ (with respect to corresponding basis, i.e., $\{dx_{j_{1}}\wedge\cdots\wedge dx_{j_{m_{i}}}\}$). One nice property is that $f_{\epsilon}$ is smooth. Another is
(6.6) ${\partial_{j}}(f_{\epsilon})=\int_{\mathbb{R}^{n}}f(y){\partial_{j}}\big(\sigma_{\epsilon}(y-x)\big)dy=\int_{\mathbb{R}^{n}}(\partial_{j}f)(y)\sigma_{\epsilon}(y-x)dy=({\partial_{j}}f)_{\epsilon}$
whenever $f$ is a smooth function. Actually ${\partial_{j}}(f_{\epsilon})=({\partial_{j}}f)_{\epsilon}$ holds for Lipschitz $f$.
Note that each $f_{i,J}$ is bounded, smooth away from the origin and may not continuous at the origin. But this does not cause problem for ${\partial_{j}}(f_{i,J})_{\epsilon}=({\partial_{j}}f_{i,J})_{\epsilon}$ because the integration by parts only needs to break into two pieces along the $x_{j}$-axis and by the boundedness of $f_{i,J}$ the contribution of integral over the $x_{j}$-axis can be ignored in (6.6) as $n>1$. Hence, each $\phi^{\epsilon}_{i}:=(d\omega_{i})_{\epsilon}=d(\omega_{i})_{\epsilon}$ is closed in $\mathbb{R}^{N_{i}+1}$ and so is their product $\phi^{\epsilon}_{1}\wedge\phi^{\epsilon}_{2}$ in $\mathbb{R}^{N_{1}+1}\oplus\mathbb{R}^{N_{2}+1}$. Besides, easy to see $\|\phi^{\epsilon}_{i}\|^{*}\leq 1$.
As argued in (6.5) with $\lim_{\epsilon\downarrow 0}\max\|\phi^{\epsilon}_{1}\wedge\phi^{\epsilon}_{2}\|^{*}=1$, the family $\phi^{\epsilon}_{1}\wedge\phi^{\epsilon}_{2}$ can serve as a calibration and thus $C_{1}\times C_{2}$ is mass-minimizing in $\mathbb{R}^{N_{1}+1}\oplus\mathbb{R}^{N_{2}+1}$. ∎
Remark 6.6.
An alternative way is to mollify $\phi=\phi_{1}\wedge\phi_{2}$ in $\mathbb{R}^{N_{1}+1}\oplus\mathbb{R}^{N_{2}+1}$. By the simpleness of $\phi_{1}$ and $\phi_{2}$, clearly $\phi$ is a smooth simple calibration away from $\big(\{0\}\times\mathbb{R}^{N_{2}+1}\big)\cup\big(\mathbb{R}^{N_{1}+1}\times\{0\}\big)$. As argued in the above proof each $\phi_{\epsilon}$ is a smooth calibration with comass no larger than one. Since the integral of $\phi_{\epsilon}$ on the truncated part of $C_{1}\times C_{2}$ within unit ball centered at the origin limits to its mass, the conclusion of Corollary 6.5 follows. Moreover, this result holds for Cartesian products of multiple Lawlor cones.
Remark 6.7.
The proof of Corollary 6.5 can work with Lawlor’s orginal non-continuous calibrations (away from the origin their potentials are Lipschitz forms). However, not like the study in [z] where scope can be completely reduced to local, here in order to establish Lemma 6.4 we need to suitably extend local calibration to a global one for our purpose and smooth calibration away from the singular set is particularly preferred.
Remark 6.8.
Given two mass minimizing currents or cones, whether their Cartesian product is minimizing is still largely open. Special cases got verified in [M].
@@
We can repeat the argument before Lemma 6.3 for contradiction. Fix $q\in M_{1}$ and $p=p_{1}$ or $p_{2}$ one singular point of $S_{2}$ and set $x=(q,p)$. Let $\xi$ be the oriented unit top vector of $T_{q}M_{1}$. Then $\xi\lrcorner\varpi_{x}$ is a calibration form with constant coefficients in $\left(T_{x}(X_{1}\times X_{2}),(\tilde{h}_{1}\oplus\tilde{h}_{2})_{x}\right)$. Since $\varpi$ calibrates $M_{1}\times S_{2}$, it can be seen that the restriction of $\xi\lrcorner\varpi_{x}$ to $\{0\}\times\mathbb{R}^{N_{2}+1}$ calibrates the tangent cone of $S_{2}$ at $p$. This contradicts with our construction which is based on local model from Theorem 1.4 - the tangent cone of $S_{2}$ at each singular point cannot support any calibration with constant coefficients.